Brian visits game-on. He is interested in playing two games, that is rocket league and horizon chase. It costs r5 to play rocket league and r2 to play horizon chase. Brian has r20 to use in playing games. Will Brian be able to play 3 games each of rocket league and horizon chase?


1
Expert's answer
2020-08-31T18:53:30-0400

Here given that Brian interesting in play two games that is rocket league and horizon chase.


rocket league has cost of playing = r5

so, cost for playing 3 game of rocket league = 3*5 = r15


horizon chase has cost of playing = r2

so, cost for playing 3 game of horizon chase = 3*2 = r6


so total cost for playing 3 of each game = r15 + r6 = r21

but , Brian has r20 to use in playing games which is less than the total cost for playing 3 of each game. so, Brian will be NOT ABLE to play 3 games each of rocket league and horizon chase.


ANSWER :

Brian will be NOT ABLE to play 3 games each of rocket league and horizon chase